# Modeling complex particle suspensions: perspectives on the rigid multiblob method

### v1

@@ I Introduction
Suspensions of small, micron-sized, particles are ubiquitous in biological, environmental and industrial systems. Their typical size and speed are such that, at their scale, the Reynolds number is very small (Re $\ll 1$) and the surrounding flow follows the Stokes equations. In this limit, the hydrodynamic interactions between the particles are long-ranged, meaning that all particle motions are instantaneously coupled through the fluid.
@@
The interplay between the individual behavior of complex particles and their suspension dynamics has attracted significant attention in the past decades. Experimental methods for synthesizing complex suspensions and numerical tools for modeling them have developed rapidly. However, significant hurdles remain in the numerical domain. Indeed, the modeling of complex suspensions poses multiple challenges due to particle geometry, thermal fluctuations with multiplicative noise, nontrivial boundary conditions, kinematic constraints, active deformations or multiphysics couplings (see Fig.1). In addition, these effects must be accurately captured to bridge the gap between the particle scale and the macroscopic suspension scale. Traditional numerical methods, such as Boundary Element Methods, are accurate but cannot handle thermal fluctuations and are restricted to few particles of anisotropic shape due to their high computational cost.
@@
An alternative approach is the rigid multiblob framework. In the rigid multiblob method (RMB), also called composite bead model [9], rigid bodies are discretized using a collection of minimally resolved spherical markers, also called blobs, or beads, which are constrained to move as a rigid body. This idea first emerged nearly 50 years ago [10] to compute the friction coefficients of non-spherical aggregates. Nowadays, the multiblob method is still widely used to compute hydrodynamic coefficients (mobility, diffusion and resistance matrices) of rigid bodies such as colloids, molecules or clusters [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 9]. However, since the 1980s, and more particularly in the past decade, several groups across the world have extensively developed the method to improve its performances and extend its domains of application [28, 29, 30, 31, 32, 9, 33, 34, 35, 36, 37, 38, 39, 40].
@@
In this Perspective, we present recent efforts to handle large numbers of particles of arbitrary shapes and multiple couplings simultaneously with a high and controllable accuracy at low computational cost. These recent developments of the rigid multiblob method open new perspectives for the modeling of complex particle suspensions and their intricate phenomena. We first introduce the governing equations of complex suspension in Section II, with an emphasis on the modeling challenges that they represent. Then we introduce rigid multiblob framework in Section III and present recent developments that successfully address these modeling challenges. Finally we showcase some examples and applications where the method has been used in Section IV and outline future directions regarding the development of the method and its use to address new physical phenomena in Section V.
@@
The framework presented in this article is implemented in a code which is free software, user friendly and continuously enriched by a group of collaborators in a GitHub repository (https://github.com/stochasticHydroTools/RigidMultiblobsWall). This code has been actively used in the community, in combination with experiments and theory, to study a variety of systems involving (micro-)swimmers in various geometries [41, 36, 42, 39, 42], passive, active and reactive colloids [43, 44, 45, 46, 47, 38], diffusing molecules [48, 49], suspensions of sedimenting cubes [50], viruses [51], functionalized nanoparticles [52], Möbius strips [53], branched particles suspensions [37, 54], fibers in structured environments [55, 56], coated drops [57], magnetic spheres [58], chains [59] or sheets [60].
@@ II.1 Governing equations for the hydrodynamic problem
where $\bm{\lambda}(\bm{r})$ is a constraint force that enforces the boundary condition (3). In absence of surface velocity ($\bm{u}_{s}=\bm{0}$), $-\bm{\lambda}$ is the hydrodynamic traction exerted on the bodies by the fluid, i.e. $-\bm{\lambda}(\bm{r})=\bm{\sigma}\cdot\bm{n}$ where $\bm{\sigma}=-p\bm{I}+2\eta(\bm{\nabla}\bm{v}+\bm{\nabla}\bm{v}^{T})$ is the fluid stress tensor. The adjoint of the geometric operator $\bm{\mathcal{K}}$ can be used to write the balance of force and torque for all bodies as
@@ II.1.1 The mobility problem
where $\bm{N}=\bm{N}(\bm{x})$ is the $6M\times 6M$ body mobility matrix that couple the forces and torques acting on the rigid bodies with their velocities, and $\widetilde{\bm{N}}$ is a linear operator ($\widetilde{\bm{N}}:L^{2}(\partial\mathcal{B},\mathbb{R}^{3})\rightarrow\mathbb{R}^{6M}$)111 $L^{2}(\partial\mathcal{B},\mathbb{R}^{3})$ is the space of vector-valued functions in $\mathbb{R}^{3}$, i.e. $\bm{u}_{s}$, defined on the manifold $\partial\mathcal{B}$ (the particle surface) whose euclidean norm is square integrable on $\partial\mathcal{B}$, i.e. $\int_{\partial\mathcal{B}}\|\bm{u}_{s}\|_{2}^{2}\mathrm{d}S<\infty$. We think this function space is a safe choice for $\bm{u}_{s}$, but it might not be the most appropriate., where $\partial\mathcal{B}$ is the particle surface, which we will call “surface mobility operator”, mapping the surface motion to the particle velocities. As shown by Swan et al. [31], which generalizes the work of Stone and Samuel [66], $\widetilde{\bm{N}}$ can be obtained for a single particle using the Lorentz reciprocal theorem.
@@ II.2 Accounting for thermal fluctuations
If the suspended particles are small enough, they will also move randomly due to thermal fluctuations in the surrounding fluid. In the overdamped limit, the effects of Brownian motion are captured through the inclusion of random velocities, turning the equations of motion into a system of stochastic differential equations given by the fluctuating Stokes equations
@@
$\nabla\cdot{\bm{v}}$ $=0,$ (10)
where the right hand side includes the divergence of a stochastic stress tensor $\bm{\mathcal{Z}}$ responsible for the thermal fluctuations [68, 69]. The stochastic stress tensor is delta correlated in space and time
$\langle\mathcal{Z}_{ij}(\bm{x},t)\rangle$ $=0,$ (11)
@@
[…]
@@
The first is the random particle velocity increments due thermal agitation given by $\bm{U}_{\text{th}}=\sqrt{2k_{B}T}\bm{N}^{1/2}\bm{Z}$, with $\bm{N}^{1/2}$ being the square root of the mobility matrix, and $\bm{Z}$ a vector of independent white noise processes. The dependence on the mobility matrix ensures that the fluctuation-dissipation theorem is satisfied, a necessary condition for the Boltzmann distribution to be recovered at equilibrium.
Along with the random velocities, the overdamped equations of motion require the inclusion of the thermal drift term, $k_{B}T\partial_{\bm{x}}\cdot\bm{N}$, which is the divergence of the mobility matrix with respect to the particle positions and orientations [73]. This ensures that the stochastic differential equation with an Itô interpretation of the stochastic integral yields dynamics consistent with those of Smoluchowski’s equation for the corresponding probability distribution. Neglecting this term in Brownian Dynamics simulations, which is sometimes the case in the literature, leads to erroneous equilibrium particle distributions [74, 75] and biases the dynamics of out-of-equilibrium systems [35, 76].
inline,backgroundcolor=blue!02inline,backgroundcolor=blue!02todo: inline,backgroundcolor=blue!02 Modelling challenges:
Computing the square root of the mobility matrix $\bm{N}^{1/2}$ with classical linear algebra, such as Cholesky decomposition, is costly as it scales badly with the particle number $O(M^{3})$. Though several methods have been introduced to accelerate the matrix square root computation [77, 78, 79, 80], the inclusion of thermal fluctuations has often limited simulations to having very small particle numbers or ignoring completely hydrodynamic interactions between the particles. The additional drift term $k_{B}T\partial_{\bm{x}}\cdot\bm{N}$ requires computing the divergence of the mobility matrix with respect to all the particle positions and orientations, which incurs a high computational cost and cannot be done analytically for complex particle shapes. In the past decade, significant advances have been made to compute these terms at a nearly linear computational cost ($O(M)$) in suspensions of spherical particles [78, 79, 74, 81, 75], but their generalization to nonspherical shapes has been lacking, and traditional methods for more complex shapes, such as the Boundary Element Method [67] or the Method of Regularized Stokeslet [82], cannot handle thermal fluctuations.
@@ II.4 Additional couplings: example with reactive particles
[…]
$D\bm{\nabla}^{2}c=0,$ (19)
with diffusion coefficient $D$.
@@ III.1 Rigid multiblob model for the Stokes problem
In the rigid multiblob method (RMB) rigid bodies are discretized using a collection of minimally resolved spherical nodes of hydrodynamic radius $a$, also called blobs, which are constrained to move as a rigid body. Figure 3 shows examples of various geometries discretized with the RMB.
@@
where $\bm{r}_{i}$ is the position of blob $i$ belonging to body $p$ ($\mathcal{B}_{p}$). The boundary condition on the particle surfaces (3), as in collocation methods [67], is evaluated at each blob
@@
where $\bm{v}(\bm{r}_{i})$ is the fluid velocity at the position of blob $i$, $\bm{M}$ is the $3N_{b}\times 3N_{b}$ mobility matrix that mediates the hydrodynamic interactions between blobs, and $N_{b}$ is the total number of blobs in the system. The term $\bm{M}_{ij}$ couples the force acting on the blob $j$ to the velocity of the blob $i$. Depending on the geometry of the domain under consideration, this matrix can be written explicitly if the Green’s function of the problem is known, or its action can be computed with matrix-free methods relying on grid-based Stokes solvers, such as Immersed Boundary Methods [107] or the Force Coupling Method [108] (see Section III.3.2). A well-known example of regularized Green’s function is the Rotne-Prager-Yamakawa (RPY) approximation [109, 110], where the regularization is done by a double convolution of the Stokes Green’s function, $\bm{\mathcal{G}}(\bm{r},\bm{r}^{\prime})$, with Dirac delta functions defined on the surface of a sphere of radius $a$:
@@
The well-known RPY mobility matrix was originally developed to provide a symmetric positive definite, pairwise approximation of the mobility matrix for a collection of spherical particles of equal radii in an unbounded domain [110]. Extensions of the RPY matrix for particles of different radii [111], in a background shear flow [110], above a fluid-fluid interface [112] or a no-slip boundary [113] with arbitrary periodicities [114, 115], or in a spherical cavity [116] are available in the literature. Section III.3.1 enumerates fast methods to compute the action of the RPY tensor.

### v2

@@ I Introduction
Suspensions of small, micron-sized, particles are ubiquitous in biological, environmental and industrial systems. Their typical size and speed are such that, at their scale, the Reynolds number is very small (Re 1$\ll$ 1≪ 1) and the surrounding flow follows the Stokes equations. In this limit, the hydrodynamic interactions between the particles are long-ranged, meaning that all particle motions are instantaneously coupled through the fluid.
@@
The interplay between the individual behavior of complex particles and their suspension dynamics has attracted significant attention in the past decades. Experimental methods for synthesizing complex suspensions and numerical tools for modeling them have developed rapidly. However, significant hurdles remain in the numerical domain. Indeed, the modeling of complex suspensions poses multiple challenges due to particle geometry, thermal fluctuations with multiplicative noise, nontrivial boundary conditions, kinematic constraints, active deformations or multiphysics couplings (see Fig.1). In addition, these effects must be accurately captured to bridge the gap between the particle scale and the macroscopic suspension scale. Traditional numerical methods, such as Boundary Element Methods, are accurate and can now handle large collections of spherical [9], or even spheroidal particles [10], but they remain restricted to few particles of complex shape due to their high computational cost and cannot handle thermal fluctuations (see Ref. [11] for an exception in 2D).
@@
An alternative approach is the rigid multiblob framework. In the rigid multiblob (RMB) method, also called composite bead model [12], or Stokesian Dynamics of rigid assemblies [13], rigid bodies are discretized using a collection of minimally resolved spherical markers, also called blobs, or beads, which are constrained to move as a rigid body. This idea first emerged nearly 50 years ago [14] to compute the friction coefficients of non-spherical aggregates. Unlike point-forces, i.e. Stokeslets, which are singular, these blobs have a finite size which allows to compute their hydrodynamic interactions, and thus the hydrodynamic coefficients of the aggregates, with regularized Green’s functions, i.e. mobility matrices, which were first derived by Rotne &\&& Prager [15] and Yamakawa [16] in the early 70s. This seminal multiblob idea was further formalized and developed in the late 1970s and early 1980s [17, 18, 19], see [20] for a review of these pioneering works. Nowadays, the multiblob method is still widely used to compute hydrodynamic coefficients (mobility, diffusion and resistance matrices) of rigid bodies such as colloids, molecules or clusters [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 12]. However, since the 1990s, and more particularly in the past decade, several groups across the world have extensively developed the method to extend its domains of application beyond the motion of passive colloidal clusters, but also to improve its numerical performances, using efficient summation techniques, modern linear algebra algorithms and grid-based numerical methods [38, 39, 40, 12, 41, 42, 43, 44, 45, 46, 47, 48].
While it shares some similarities with the well-known method of Regularized Stokeslets [49, 35], in which bodies are discretized with regularized point forces, the rigid multiblob method additionally handles thermal fluctuations and effectively couples with grid-based methods (such as Immersed Boundary methods [50, 51]) to solve hydrodynamic interactions between blobs, whereas the former can only be used for deterministic problems and relies exclusively on analytical Green’s functions.
@@
In this Perspective, intended for a wide audience of fluid dynamicists and soft matter physicists, we present recent developments of the rigid multiblob method to model suspensions containing a large number of particles of arbitrary shapes and subject to multiple physical couplings, with controllable accuracy and high scalability. These recent efforts open new perspectives for the modeling of complex particle suspensions and their intricate phenomena. We first introduce the governing equations of complex particle suspensions in Section II, with an emphasis on the modeling challenges that they represent. Then we introduce the rigid multiblob framework in Section III and present recent developments that successfully address these modeling challenges. Finally we showcase some examples and applications where the method has been used in Section IV and outline future directions regarding the development of the method and its use to address new physical phenomena in Section V.
@@
The framework presented in this article is implemented in a code which is free software, user friendly for non-experts and continuously enriched by a group of collaborators in a GitHub repository (https://github.com/stochasticHydroTools/RigidMultiblobsWall). The repository contains many hands-on examples that users can run on a basic laptop. This code has been actively used in the community, in combination with experiments and theory, to study a variety of systems involving (micro-)swimmers in various geometries [52, 44, 53, 47, 53], passive, active and reactive colloids [54, 55, 56, 57, 58, 59, 46, 60], diffusing molecules [61, 62], suspensions of sedimenting cubes [63], viruses [64], functionalized nanoparticles [65], Möbius strips [66], branched particles suspensions [45, 67], fibers in structured environments [68, 69], coated drops [70], magnetic spheres [71], chains [72] or sheets [73].
@@ II.1 Governing equations for the hydrodynamic problem
where $\bm{\lambda}(\bm{r})$ is a constraint force per unit area that enforces the boundary condition (3). In absence of surface velocity ($\bm{u}_{s}=\bm{0}$), $-\bm{\lambda}$ is the hydrodynamic traction exerted on the bodies by the fluid, i.e. $-\bm{\lambda}(\bm{r})=\bm{\sigma}\cdot\bm{n}$ where $\bm{\sigma}=-p\bm{I}+2\eta(\bm{\nabla}\bm{v}+\bm{\nabla}\bm{v}^{T})$ is the fluid stress tensor [78, 79]. The adjoint of the geometric operator $\bm{\mathcal{K}}$ can be used to write the balance of force and torque for all bodies as 111 The adjoint property ensures that the power dissipated by the motion of the particle surface integrated over the whole surface matches the power dissipated by the motion of a rigid particle. For a particle without surface velocity ($\bm{u}_{s}=\bm{0}$) $P$ $=\int_{\partial\mathcal{B}_{p}}\bm{\lambda}^{T}(\bm{r})\,\bm{v}(\bm{r})\;\mathrm{d}S_{\bm{r}}=\int_{\partial\mathcal{B}_{p}}\bm{\lambda}^{T}(\bm{r})\,(\bm{\mathcal{K}}\bm{U})(\bm{r})\;\mathrm{d}S_{\bm{r}}=\bm{U}^{T}\bm{\mathcal{K}}^{T}\bm{\lambda}=\bm{U}^{T}\bm{F}=\bm{u}\cdot\bm{f}+\bm{\tau}\cdot\bm{\omega}.$ (6)
@@ II.1.1 The mobility problem
where $\bm{N}=\bm{N}(\bm{x})$ is the $6M\times 6M$ body mobility matrix that couple the forces and torques acting on the rigid bodies with their velocities, and $\widetilde{\bm{N}}$ is a linear operator ($\widetilde{\bm{N}}:L^{2}(\partial\mathcal{B},\mathbb{R}^{3})\rightarrow\mathbb{R}^{6M}$)222 $L^{2}(\partial\mathcal{B},\mathbb{R}^{3})$ is the space of vector-valued functions in $\mathbb{R}^{3}$, i.e. $\bm{u}_{s}$, defined on the manifold $\partial\mathcal{B}$ (the particle surface) whose euclidean norm is square integrable on $\partial\mathcal{B}$, i.e. $\int_{\partial\mathcal{B}}\|\bm{u}_{s}\|_{2}^{2}\mathrm{d}S<\infty$. We think this function space is a safe choice for $\bm{u}_{s}$, but it might not be the most appropriate., where $\partial\mathcal{B}$ is the particle surface, which we will call “surface mobility operator”, mapping the surface motion to the particle velocities. As shown by Swan et al. [39], which generalizes the work of Stone and Samuel [82], $\widetilde{\bm{N}}$ can be obtained for a single particle using the Lorentz reciprocal theorem.
@@ II.2 Accounting for thermal fluctuations
If the suspended particles are small enough, they will also move randomly due to thermal fluctuations in the surrounding fluid. In order to satisfy the correct diffusion in the suspension, the statistics of the resulting particle fluctuating velocities, $\bm{U}_{\text{th}}$, must satisfy the fluctuation-dissipation theorem [83, 84]
$\langle\bm{U}_{\text{th}}\rangle$ $=\bm{0},$ (10)
$\langle\bm{U}_{\text{th}}(t)\bm{U}_{\text{th}}(t^{\prime})^{T}\rangle$ $=\bm{D}\delta(t-t^{\prime})=2k_{B}T\bm{N}\delta(t-t^{\prime}),$ (11)
[…]
In order to incorporate the effects of Brownian motion in the fluid, a stochastic stress tensor $\bm{\mathcal{Z}}$ [87, 88] is added into the right hand side of the Stokes equations
@@
$\nabla\cdot{\bm{v}}$ =0=0,= 0 , (13)
The tensor $\bm{\mathcal{Z}}$ encodes the stress generated by the random motion and collisions of the fluid molecules, which, at scales much larger than the molecular scale, can be modeled as delta correlated in space and time
$\langle\mathcal{Z}_{ij}(\bm{x},t)\rangle$ =0=0,= 0 , (14)
@@
The ensemble average of $\bm{\mathcal{Z}}$ is zero because $\bm{\mathcal{Z}}$ represents the random fluctuations around the mean deterministic flow [87, 85]. The magnitude of the correlations, proportional to the thermal energy $k_{B}T$, and their symmetries ensure that the fluctuation-dissipation theorem (11) is satisfied by Eqs. (12)-(15) [87, 88].
@@
The first is the random particle velocity increments due thermal agitation given by $\bm{U}_{\text{th}}=\sqrt{2k_{B}T}\bm{N}^{1/2}\bm{Z}$, with $\bm{N}^{1/2}$ being the square root of the mobility matrix (i.e. $\bm{N}^{1/2}\left(\bm{N}^{1/2}\right)^{T}=\bm{N}$), and $\bm{Z}$ a vector of independent white noise processes. The noise dependence on the mobility matrix ensures that the fluctuation-dissipation theorem (11) is satisfied, a necessary condition for the Boltzmann distribution to be recovered at equilibrium [92]. Note that while both $\bm{\mathcal{Z}}$ and $\bm{Z}$ present in Eqs. (12) and (16) represent white noise terms they are not equivalent because the first is a continuous field defined on the fluid while the second is a discrete vector on the particles.
Along with the random velocities, the overdamped equations of motion require the inclusion of the thermal drift term, $k_{B}T\partial_{\bm{x}}\cdot\bm{N}$, which is the divergence of the mobility matrix with respect to the particle positions and orientations [93]. This ensures that the stochastic differential equation with an Itô interpretation of the stochastic integral yields dynamics consistent with those of Smoluchowski’s equation for the corresponding probability distribution.
[…]
Computing the square root of the mobility matrix $\bm{N}^{1/2}$ with classical linear algebra, such as Cholesky decomposition, is costly as it scales badly with the particle number $O(M^{3})$. Though several methods have been introduced to accelerate the matrix square root computation [98, 94, 99, 100], the inclusion of thermal fluctuations has often limited simulations to having very small particle numbers or ignoring completely hydrodynamic interactions between the particles. The additional drift term $k_{B}T\partial_{\bm{x}}\cdot\bm{N}$ requires computing the divergence of the mobility matrix with respect to all the particle positions and orientations, which incurs a high computational cost and cannot be done analytically for complex particle shapes. In the past decade, significant advances have been made to compute these terms at a nearly linear computational cost ($O(M)$) in suspensions of spherical particles [94, 99, 92, 101, 95], but their generalization to nonspherical shapes has been lacking, and traditional methods for more complex shapes, such as the Boundary Element Method [78] or the Method of Regularized Stokesleta [49], cannot handle thermal fluctuations.
@@ II.4 Additional couplings: example with reactive particles
Assuming that the magnitude of the phoretic flows is small and the solute diffusivity $D$ is large, the Péclet number characterizing the motion of the solute particles is small enough to neglect solute advection. Thus, in the fluid domain the concentration of solute $c=c(\bm{r})$ diffuses according to the Laplace equation
$D\bm{\nabla}^{2}c=0.$ (22)
@@ III.1 Rigid multiblob model for the Stokes problem
In the rigid multiblob (RMB) method rigid bodies are discretized using a collection of minimally resolved spherical nodes of hydrodynamic radius $a$, also called blobs, which are constrained to move as a rigid body. Figure 3 shows examples of various geometries discretized with the RMB.
@@
where $\bm{r}_{i}$ is the position of blob $i$ belonging to body $p$ ($\mathcal{B}_{p}$). Note that here $\bm{\lambda}_{i}$ now represents finite forces and not density forces as in the continuum formulation since it includes the quadrature weights of the blobs. The boundary condition on the particle surfaces (3), as in collocation methods [78], is evaluated at each blob
@@
where $\bm{v}(\bm{r}_{i})$ is the fluid velocity at the position of blob $i$, $\bm{M}$ is the $3N_{b}\times 3N_{b}$ mobility matrix that mediates the hydrodynamic interactions between blobs, and $N_{b}$ is the total number of blobs in the system. The term $\bm{M}_{ij}$ couples the force acting on the blob $j$ to the velocity of the blob $i$. Depending on the geometry of the domain under consideration, this matrix can be written explicitly if the Green’s function of the problem is known, or its action can be computed with matrix-free methods relying on grid-based Stokes solvers, such as Immersed Boundary Methods [50] or the Force Coupling Method [51] (see Section III.3.2).
A well-known example of regularized Green’s function is the Rotne-Prager-Yamakawa (RPY) approximation [15, 126], where the regularization is done by a double convolution of the Stokes Green’s function, $\bm{\mathcal{G}}(\bm{r},\bm{r}^{\prime})$, with Dirac delta functions defined on the surface of a sphere of radius $a$:
@@
Wajnryb et al. showed that, for non-overlapping blobs, an equivalent definition of the RPY mobility is [126]
$\bm{M}_{ij}=\bm{M}(\bm{r}_{i},\bm{r}_{j})=\left[\bm{I}+\frac{a^{2}}{6}\bm{\nabla}^{2}_{\bm{r}^{\prime}}\right]\left[\bm{I}+\frac{a^{2}}{6}\bm{\nabla}^{2}_{\bm{r}^{\prime\prime}}\right]\bm{\mathcal{G}}(\bm{r}^{\prime},\bm{r}^{\prime\prime})|_{\bm{r}^{\prime\prime}=\bm{r}_{j}}^{\bm{r}^{\prime}=\bm{r}_{i}},$ (29)
which makes evident the role of the blob radius as a regularization parameter.
The well-known RPY mobility matrix was developed to provide a symmetric positive definite, pairwise approximation. The use of a symmetric positive definite matrix $\bm{M}$ is one of the main advantages of the RMB as it allows to incorporate Brownian fluctuations efficiently as shown in Sec. III.2. In contrast, the Regularized Stokeslets method uses a mobility $\bm{M}$ that is not positive definite by design [35] and, current Boundary Element Methods use singular and near singular quadrature rules to compute the action of the Stokes Green’s function, $\bm{\mathcal{G}}(\bm{r},\bm{r}^{\prime})$ that generate non-symmetric mobility matrices $\bm{M}$ and thus, cannot incorporate Brownian fluctuations efficiently (see Ref. [11] for an exception in 2D). As explained in Sec. III.3.2, the mobility matrix of grid-based- methods is symmetric positive definite by design and has the same structure (29) as the RPY mobility to order $\mathcal{O}\left(a^{4}\right)$.
Since $\bm{M}$ is positive definite it is not necessary to employ special quadrature rules to evaluate near or self interactions. However, the formula (28) introduces a regularization error proportional to the blob radius [46]. As a typical discretization of a rigid body surface employs $N_{b}\sim 1/a^{2}$ blobs, the convergence rate scales with the number of blobs as $\sim N_{b}^{-1/2}$. Despite this slow convergence it is possible to use optimized grids to obtain results accurate up to a few percent as far as blobs of different bodies do not overlap [127, 46]. We illustrate the accuracy of the RMB method in a couple of examples in Secs. III.4-III.5 and discuss optimized grids in Sec. V.
The original RPY mobility computes hydrodynamic interactions between spherical particles of equal radii in an unbounded domain [126]. Extensions of the RPY matrix for particles of different radii [128], in a background shear flow [126], above a fluid-fluid interface [129] or a no-slip boundary [130] with arbitrary periodicities [131, 132], or in a spherical cavity [133] are available in the literature. Section III.3.1 enumerates fast methods to compute the action of the RPY tensor.